# Global Deep Forecasting with Patient-Specific Pharmacokinetics

### v8

@@ Abstract
Forecasting healthcare time series data is vital for early detection of adverse outcomes and patient monitoring. However, it can be challenging in practice due to variable medication administration and unique pharmacokinetic (PK) properties of each patient. To address these challenges, we propose a novel hybrid global-local architecture and a PK encoder that informs deep learning models of patient-specific treatment effects. We showcase the efficacy of our approach in achieving significant accuracy gains in a blood glucose forecasting task using both realistically simulated and real-world data. Our PK encoder surpasses baselines by up to 16.4% on simulated data and 4.9% on real-world data for individual patients during critical events of severely high and low glucose levels. Furthermore, our proposed hybrid global-local architecture outperforms patient-specific PK models by 15.8%, on average. Our code is available at https://github.com/PotosnakW/neuralforecast/tree/pk_paper_code
@@ 1 Introduction
Leveraging the predictive power of exogenous variables presents unique challenges, especially in healthcare, where the frequency of patient signals and exogenous factors, such as medication doses or treatments, often exhibit a mismatch in temporal resolution. Furthermore, while the drug dose can be observed and recorded, the body’s interaction with the drug cannot be directly observed and is typically inferred from its plasma concentration, as illustrated in Fig. 1. Consequently, the recorded drug dose alone cannot capture time-dependent drug kinetics, including processes such as absorption and elimination.
@@
Consider the task of blood glucose monitoring for patients with diabetes who must regularly monitor and regulate their blood glucose level. Careful and prompt administration of insulin and intake of carbohydrates is required to prevent hyper- and hypoglycemic events, which can cause irreversible organ damage or, in severe cases, be fatal. Hyperglycemic and hypoglycemic episodes are responsible for numerous hospital admissions each year. A 2009 report on hospital admissions indicates over 500,000 patient-days and costs of \$2.4 billion each year for hyperglycemia alone (Kitabchi et al., 2009). Another report on hypoglycemic hospitalizations suggests a stable rate of 5.6-7.3 admissions per 1,000 patients annually between 2016-2019 (Pilla et al., 2021). With over 38 million patients with Diabetes Mellitus in the United States alone, this equates to a conservative estimate of 212,800 patients hospitalized for hypoglycemic episodes each year.
For healthcare challenges like this, advanced forecasting technologies can help patients and clinicians monitor, anticipate, and proactively address emerging abnormalities to prevent adverse events. However, accurate forecasting requires models that can leverage exogenous information, such as treatments, while capturing latent treatment kinetics and their effects to generate reliable forecast trajectories. To address this, we propose a hybrid model that integrates cohort-level information with patient-specific pharmacokinetic characteristics to improve the predictability of blood glucose responses to treatment.
@@
    Hybrid Global-Local Architecture. We bolster global architectures (with parameters shared across patients) with a learnable patient-specific encoder.
@@ 1.1 Pharmacokinetics Background
PK modeling is a natural solution to the aforementioned problem, as it can reveal the effects of the medication over time through knowledge of drug movement within the body. For example, PK provides a well-established understanding of the changes in plasma insulin concentration over time based on drug absorption and elimination characteristics  (Binder et al., 1984). Empirical and invasive studies show that concentration-time profiles for subcutaneous injections follow right-skewed curves. The extent of skew varies with insulin type and patient-specific factors, as shown in Fig. 2, where the area under the concentration curve measures drug bioavailability, or the fraction of the absorbed dose that reaches the intended site of systemic circulation intact. Most drugs used in clinical practice are assumed to follow linear PK (Rosenbaum, 2017). Under this model, the area under the concentration curve is directly proportional to the drug dose. A complete reference for supporting PK information regarding concentration-time profiles, insulin analogs, bioavailability, first-order kinetics, and linear PK can be found in Appendix B.
Figure 2: Pharmacokinetic modeling of plasma insulin concentration over time for various insulin types.
@@ 2 Related Work
Prior works have looked at modeling treatment effects on blood glucose levels. Miller et al. introduced one of the first hybrid statistical-physiological models, the Deep T1D Simulator (DTD-Sim), which integrates glucose-insulin dynamics with deep learning through a deep state-space model (Miller et al., 2020). Odnoblyudova et al. improved the prediction of blood glucose using multi-output Gaussian processes (MOGP) to model the composite effects of meal components, such as carbohydrates and fats (Odnoblyudova et al., 2023). Unlike meals, insulin is a homogeneous protein hormone, making component-based modeling inapplicable to our task (Odnoblyudova et al., 2023). Instead, our objective is to capture the holistic patient-specific effects of the drug and the composite effects of multiple doses.
@@
Previous work has developed individualized networks that overlook the benefits of developing global models that learn from various, diverse data series. Such cross-learning on diverse datasets can enhance generalization abilities, as demonstrated with Time Series Foundation Models (TSFMs) (Goswami et al., 2024). Our hybrid global-local architecture enables cohort-level training to develop a global model applicable to multiple patients while also preserving patient-specific information for personalized forecasts. Regarding architectures, models such as the UVA / Padova simulator depend on predefined ODEs for glucose dynamics, restricting adaptability to other targets and exogenous inputs (Visentin et al., 2016; Miller et al., 2020). Similarly, deep learning approaches for sparse exogenous variables often rely on fixed base models like LSTMs (Rubin-Falcone et al., 2022). Our proposed contributions support diverse deep learning models, and are flexible to forecasting targets and input variables, ensuring future adaptability as more advanced tools emerge. Other prior work that proposes the integration of PK knowledge with deep learning focuses on domains such as long-term disease progression in cancer and COVID-19, where such PK models are not directly applicable to blood glucose and insulin kinetics. (Hussain et al., 2021; Qian et al., 2021). Our study provides a formal framework for integrating PK in deep learning under linear PK assumptions that are applicable to many drugs used in clinical practice (Rosenbaum, 2017).
@@ 3 Methods
We aim to forecast blood glucose levels 30 minutes into the future consistent with prior work (Rubin-Falcone et al., 2022; Xie and Wang, 2020; Li et al., 2019b), based on a 10 hour history of blood glucose, carbohydrate (CHO), insulin basal, and insulin bolus values. We choose a 10-hour duration to balance sufficient historical information with our GPU memory constraints. Moreover, the duration of rapid-acting insulin (Humalog and Novolog) used in insulin pumps by patients in our cohort typically ranges from 2 to 5 hours (Clinic, 2018; Association, n.d.). Thus, a 10-hour timespan fully captures multiple dose durations, enabling the model to learn treatment effects across various dosages and durations.
@@ 3.1 PK Encoder
Medication administrations are represented as instantaneous sparse variables in time series data. We propose a novel PK encoder that informs deep learning models of time-dependent plasma drug concentration to enable more accurate forecasting of treatment outcomes as shown in Fig. 3. The encoder consists of a function, $C$, that takes in sparse exogenous variables and generates a concentration-time profile for a specified dose, $\textbf{x}^{(i)}$, using the following equation:
$C(t,\textbf{x}^{(i)},\textbf{k}^{(i)})=\frac{\textbf{x}^{(i)}}{\textbf{k}^{(i)}\sqrt{2\pi}t}\exp\{-\frac{1}{2{(\textbf{k}^{(i)})}^{2}}(\log(t)-1)^{2}\}$ (1)
where $t$ is the time and $\textbf{k}^{(i)}$ is a PK parameter affecting elimination rate specific to the patient, $i$. Here, $C$ is equivalent to the log-normal probability density function (pdf) with $\sigma=\textbf{k}$ and $\mu=1$ and a scaling factor, x. The intuition for the log-normal shape of the concentration curves stems from PK literature, as outlined in Section 1.1 and Appendix B.
@@
Under linear PK assumptions, the PK parameters, including bioavailability, are assumed to be constant with respect to the drug. Given this assumption, we show mathematically that the area under the concentration curve is bounded by the dose in Proposition A.1 in Appendix A.4. To model bioavailability and incorporate dose information directly into the concentration curves, we use the dose x, as a scaling factor in $C$. This ensures that the area under the curve integrates to the dose value, under the simplifying assumption that all insulin administrations adhere to 100% bioavailability, which we show is upper bounded by the dose. Scaling the concentration curve changes its magnitude, but does not alter the PK parameters or shape of the curve.
Let $N$ be the number of subjects in the data. The PK parameter $\textbf{k}\in[0,1]^{N\times 1}$ that controls the shape of the concentration curve, is represented as network embedding weight that is initialized between values of 1 and 2 prior to model training. The initial k values are tuned as model hyperparameters.
The encoder is trained end-to-end with the model. During training, sparse insulin features are passed into the PK encoder along with their respective $\textbf{k}^{(i)}$ parameters, as shown in Fig. 4. The encoder then replaces sparse medication features in the training window with the corresponding concentration curves. For a single dose, x, of insulin administered at time, $t$, we obtain the concentration, $\widecheck{\textbf{x}}^{(i)}$, at that time for patient, $i$:
$\widecheck{\textbf{x}}^{(i)}$ $=C(t,\textbf{x}^{(i)},\textbf{k}^{(i)}).$ (2)
For each training step, $\textbf{k}^{(i)}$ values are updated for each patient, $i$, using gradient descent optimization.
@@
For each time step within the forecast horizon, $H$, forecasts, $\hat{\textbf{y}}$, are generated by the model, $f_{\theta}$, as a function of static features, s, treatment concentration curves, $\widecheck{\textbf{x}}$, and historical glucose values y recorded up to and including time $t$ based on the lag time, $L$,
$\hat{\textbf{y}}^{(i)}_{t+1:t+H}$ $=f_{\theta}(\textbf{y}^{(i)}_{t-L:t},\widecheck{\textbf{x}}^{(i)}_{t-L:t},\textbf{s}^{(i)}).$ (3)
@@
In the proposed context, s includes patient age, weight, one-hot encoded subject identification number, and insulin pump type. $\widecheck{\textbf{x}}_{t-L:t}$ includes concentration curves for basal insulin, $\widecheck{\textbf{x}}_{\text{basal}}$, bolus insulin, $\widecheck{\textbf{x}}_{\text{bolus}}$, and CHO, $\widecheck{\textbf{x}}_{\text{CHO}}$. We include treatment effects of carbohydrates because, similar to drugs, they undergo absorption and metabolism processes, resulting in glucose and other sugars entering the bloodstream.
@@ 3.2 Modeling Multiple Dose Effects
Multiple insulin doses may occur at close intervals in an event often referred to as “insulin stacking” as shown in Fig. 3 and Fig. 14. To model PK effects for multiple doses in an input window, we generate a concentration curve for each dose event and sum concentration curves across the windows. Under the assumption of linear pharmacokinetics (outlined in 1.1 and Appendix B), we show that the overall effect, or concentration, of $n$ doses of the same drug is equal to the sum of the individual effects which is verified mathematically in Proposition A.2 in Appendix A.4. To perform vectorized operations that enable generating concentration curves for multiple doses in an input window, the PK encoder leverages a matrix, $\textbf{w}\in\mathbb{R}^{L\times L}$, where $L$ refers to the lag time. This approach allows for efficient modeling of multiple effects in $\mathcal{O}(1)$ time. Each row of w represents a time count up to the input time where the count shifts one time step for each subsequent row. Concentration curves are generated for each row of w using the dose at the corresponding time index. If no dose is administered at a particular time, $t$, the corresponding row of w indexed by $t$ will be zeroed out. The final concentration curve feature, which can reflect the impact of multiple doses, is generated by aggregating the individual concentration curves across time steps:
w $=\begin{bmatrix}0&1&2&3&\cdots&L-1\\ 0&0&1&2&\cdots&L-2\\ \ldots&\ldots&\ldots&\ldots&\ldots&\ldots\\ 0&0&0&0&\cdots&0\end{bmatrix}\in\mathbb{R}^{L\times L}$ (4)
$\widecheck{\textbf{x}}^{(i)}$ $=\sum_{t=0}^{L}C(\textbf{w}_{t,j}\cdot f,\textbf{x}^{(i)}_{t},\textbf{k}^{(i)})$ (5)
Here, $f$ is the frequency of the data, $\textbf{x}^{(i)}\in\mathbb{R}^{L\times 1}$ a sparse exogenous time series of treatment, and $\widecheck{\textbf{x}}^{(i)}\in\mathbb{R}^{1\times L}$ is the new concentration curve feature for patient, $i$.
The PK encoder training algorithm is described in Appendix A.5, which can be integrated into any deep learning architecture capable of handling exogenous variables in addition to modeling the target variable.
Figure 4: Our hybrid global-local architecture combines global model parameters shared across patients with patient-specific pharmacokinetic (PK) parameters. Sparse treatment time series features, such as medication doses, are input into the PK encoder to generate patient-specific treatment effect curves, $\check{\textbf{x}}$. PK encoder output, $\check{\textbf{x}}$, and any other exogenous features, x, are used as deep learning model inputs. During each training step, the model learns relevant hidden states and output parameters, $\hat{\theta}$, across patients, while $\textbf{k}^{(i)}$ values are updated for each patient, $i$, using gradient descent optimization.

### v9

@@ Abstract
Forecasting healthcare time series data is vital for early detection of adverse outcomes and patient monitoring. However, it can be challenging in practice due to variable medication administration and unique pharmacokinetic (PK) properties of each patient. To address these challenges, we propose a novel hybrid global-local architecture and a PK encoder that informs deep learning models of patient-specific treatment effects. We showcase the efficacy of our approach in achieving significant accuracy gains in a blood glucose forecasting task using both realistically simulated and real-world data. Our PK encoder surpasses baselines by up to 16.4% on simulated data and 4.9% on real-world data for individual patients during critical events of severely high and low glucose levels. Furthermore, our proposed hybrid global-local architecture outperforms patient-specific PK models by 15.8%, on average.
Data and Code Availability
Institutional Review Board (IRB)
Our research does not require IRB approval.
@@ 1 Introduction
Leveraging the predictive power of exogenous variables presents unique challenges, especially in healthcare, where the frequency of patient signals and exogenous factors, such as medication doses or treatments, often exhibit a mismatch in temporal resolution. Furthermore, while the drug dose can be observed and recorded, the body’s interaction with the drug cannot be directly observed and is typically inferred from its plasma concentration, as illustrated in Fig. 1. Consequently, the recorded drug dose alone cannot capture time-dependent pharmacokinetics, including processes such as absorption and elimination.
Consider the task of blood glucose monitoring for patients with Type-1 Diabetes Mellitus (T1DM) or insulin-dependent T2DM patients who must regularly monitor and regulate their blood glucose levels. Careful and prompt administration of insulin and intake of carbohydrates is required to prevent hyper- and hypoglycemic events, which account for numerous hospital admissions annually, and can cause irreversible organ damage or, in severe cases, be fatal. A 2016 report estimated that hypoglycemia causes approximately 100,000 emergency department (ED) visits per year due to insulin-related errors, with nearly one-third resulting in hospitalization (Geller et al., 2014). Another study reports that in 2014, there were 184,255 ED visits and in-patient U.S. hospital admissions for diabetic ketoacidosis (DKA), often associated with hyperglycemia, with 70.6% occurring in T1DM patients (Benoit et al., 2020). Another study reported a trend toward increasing hospitalizations for DKA in T1DM patients from 2008 to 2018 (Kichloo et al., 2021).
For healthcare challenges like this, advanced forecasting technologies can help patients and clinicians monitor, anticipate, and proactively address emerging abnormalities to prevent adverse events. However, accurate forecasting requires models that can leverage exogenous information, such as treatments, while capturing latent treatment kinetics and their effects to generate reliable forecast trajectories. To address this, we propose a hybrid model that integrates cohort-level information with patient-specific pharmacokinetics to improve the prediction of blood glucose responses to treatment.
@@
    Hybrid Global-Local Architecture. We bolster global architectures (with parameters shared across patients) with a learnable, patient-specific encoder.
@@ 1.1 Pharmacokinetics Background
Pharmacokinetic (PK) modeling is a natural solution to the aforementioned problem, as it can reveal medication effects over time through well-established knowledge of changes in plasma insulin concentration driven by absorption and elimination characteristics  (Binder et al., 1984). Empirical and invasive studies show that concentration-time profiles for subcutaneous injections follow right-skewed curves. The extent of skew varies with insulin type and patient-specific factors, as shown in Fig. 2, where the area under the concentration curve measures drug bioavailability, or the fraction of the absorbed dose that reaches the intended site of systemic circulation intact. Most drugs used in clinical practice are assumed to follow linear PK (Rosenbaum, 2017). Under this model, the area under the concentration curve is directly proportional to the drug dose. A complete reference for supporting PK information regarding concentration-time profiles, insulin analogs, bioavailability, first-order kinetics, and linear PK can be found in Appendix B.
Figure 2: Pharmacokinetic models of plasma insulin concentration over time for various insulin types.
@@ 2 Related Work
Prior work has looked at modeling treatment effects on blood glucose levels. Miller et al. introduced one of the first hybrid statistical-physiological models, the Deep T1D Simulator (DTD-Sim), which integrates glucose-insulin dynamics with deep learning through a deep state-space model (Miller et al., 2020). Odnoblyudova et al. (2023) improved the prediction of blood glucose using multi-output Gaussian processes (MOGP) to model the composite effects of meal components, such as carbohydrates and fats. Unlike meals, insulin is a homogeneous protein hormone, making component-based modeling inapplicable to our task. Instead, our objective is to capture the holistic patient-specific effects of the drug and the composite effects of multiple doses.
@@
Previous work has focused on developing individualized networks, often overlooking the advantages of global models that learn from multiple, diverse time series—a technique known as cross-learning (Smyl, 2019; Semenoglou et al., 2021), and employed by Time Series Foundation Models (TFSMs) (Goswami et al., 2024). Our hybrid global-local architecture enables cohort-level training to develop a global model applicable to multiple patients while also preserving patient-specific information for personalized forecasts. Regarding architectures, models such as the UVA / Padova simulator depend on predefined ODEs for glucose dynamics, restricting adaptability to other targets and exogenous inputs (Visentin et al., 2016; Miller et al., 2020). Similarly, deep learning approaches for sparse exogenous variables often rely on fixed base models like LSTMs (Rubin-Falcone et al., 2022). Our proposed contributions support diverse deep learning models, and are flexible to forecasting targets and input variables, ensuring future adaptability as more advanced tools emerge. Other prior work that proposes the integration of PK knowledge with deep learning focuses on domains such as long-term disease progression in cancer and COVID-19, where such PK models are not directly applicable to blood glucose and insulin kinetics. (Hussain et al., 2021; Qian et al., 2021). Our study provides a formal framework for integrating PK in deep learning under linear PK assumptions that are applicable to many drugs used in clinical practice (Rosenbaum, 2017).
@@ 3 Methods
We aim to forecast blood glucose values 30 minutes into the future consistent with prior work (Rubin-Falcone et al., 2022; Xie and Wang, 2020; Li et al., 2019b), based on a 10-hour history of blood glucose, carbohydrate (CHO), insulin basal, and insulin bolus values. We choose a 10-hour duration to balance sufficient historical information with our GPU memory constraints. Moreover, the duration of rapid-acting insulin (Humalog and Novolog) used in insulin pumps by patients in our cohort typically ranges from 2 to 5 hours (Clinic, 2018; Association, n.d.). Thus, a 10-hour time span fully captures multiple dose durations, enabling the model to learn treatment effects across various dosages and durations. We also demonstrate that the significant performance improvement of the proposed methods holds across other various lag times in Appendix C.3.
@@ 3.1 PK Encoder
Medication administrations are recorded as instantaneous dosing events, leading to sparse observations in time series data. We propose a novel PK encoder that informs deep learning models of time-dependent plasma drug concentrations to enable more accurate forecasting of treatment outcomes as shown in Fig. 3. The encoder consists of a function $C$ that generates a concentration-time profile for a single specified dose $x$ at time $t$ using a PK parameter $k$ affecting elimination rate:
$C(t,x,k)=\frac{x}{tk\sqrt{2\pi}}\exp\{-\frac{1}{2k^{2}}(\log(t)-1)^{2}\}.$ (1)
Here, $C$ is equivalent to the log-normal probability density function with $\sigma=k$ and $\mu=1$ and a scaling factor $x$. The intuition for the log-normal shape of the concentration curves stems from PK literature, as outlined in Section 1.1 and Appendix B.
@@
Under linear PK assumptions, the PK parameters, including bioavailability, are assumed to be constant with respect to the drug. Given this assumption, we show mathematically that the area under the concentration curve is bounded by the dose in Proposition A.1 in Appendix A.5. To model bioavailability and incorporate dose information directly into the concentration curves, we use the dose $x$ as a scaling factor in $C$. This ensures that, under the simplifying assumption of 100% bioavailability, the area under the curve integrates to the dose. Scaling the concentration curve in this way affects its magnitude but does not alter PK parameters that determine its shape.
Let $N$ be the number of subjects in the dataset. The PK parameter vector $\mathbf{k}\in\mathbb{R}^{N\times 1}$ represents patient-specific embedding weights that control the shape of the patient-specific concentration curves, where $\mathbf{k}^{(i)}$ corresponds to the PK parameter $k$ for the $i$-th patient. The vector $\mathbf{k}$ is initialized with values between 1 and 2 prior to model training, with the initial value treated as a tunable hyperparameter.
The encoder is trained end to end with the model. During training, a sparse treatment feature $\mathbf{x}^{(i)}\in\mathbb{R}^{1\times L}$, representing dose administrations within an input window of length $L$ up to and including the current time $T$, is passed to the PK encoder along with the corresponding parameter $\mathbf{k}^{(i)}$ to generate a concentration curve feature $\widecheck{\mathbf{x}}^{(i)}\in\mathbb{R}^{1\times L}$, as shown in Fig. 4. In the single-dose setting, where a dose is recorded at time step $t_{d}\in\{T-L+1,\dots,T\}$, the encoder replaces the sparse treatment feature $\mathbf{x}_{T-L+1:T}^{(i)}$ with the corresponding concentration feature $\widecheck{\mathbf{x}}_{T-L+1:T}^{(i)}$, computed as:
$\widecheck{\mathbf{x}}^{(i)}_{t}$ $=\begin{cases}0,&\!\text{for }t\in[T-L+1,t_{d})\\ C\left(\nu(t-t_{d}),\mathbf{x}^{(i)}_{t_{d}},\mathbf{k}^{(i)}\right),&\!\text{for }t\in[t_{d},T]\end{cases}$ (2)
where $t$ indexes time steps within the window, $\nu$ is the sampling interval of the data, and $C(\cdot)$ computes the concentration at elapsed time $\nu(t-t_{d})$ since the dose was administered. During training, the parameter $\mathbf{k}^{(i)}$ is updated for each patient $i$ via gradient descent.
@@
Forecasts $\hat{\textbf{y}}$ are generated by the model $f_{\theta}$ for each time step within the forecast horizon $H$ as a function of historical glucose values y, treatment concentration curves $\widecheck{\textbf{x}}$, and static features s, recorded within the input window up to and including time $T$,
$\hat{\textbf{y}}^{(i)}_{T+1:T+H}$ $=f_{\theta}(\textbf{y}^{(i)}_{T-L+1:T},\widecheck{\textbf{x}}^{(i)}_{T-L+1:T},\textbf{s}^{(i)}).$ (3)
@@
In the proposed context, s includes patient age, weight, one-hot encoded subject identification number, and insulin pump type. $\widecheck{\textbf{x}}_{T-L+1:T}$ includes concentration curve values for basal insulin $\widecheck{\textbf{x}}_{\text{basal}}$, bolus insulin $\widecheck{\textbf{x}}_{\text{bolus}}$, and CHO $\widecheck{\textbf{x}}_{\text{CHO}}$. We include treatment effects of carbohydrates because, similar to drugs, they undergo absorption and metabolism processes, resulting in glucose and other sugars entering the bloodstream.
@@ 3.2 Modeling Multiple Dose Effects
Multiple insulin doses may occur at close intervals in an event often referred to as “insulin stacking” as shown in Fig. 3 and Fig. 15. To model PK effects for multiple doses within an input window, we generate a concentration curve for each dose event and sum the concentration curves within the window. Under the assumption of linear PK (outlined in 1.1 and Appendix B), we show that the overall effect, or concentration, of $n$ doses of the same drug is equal to the sum of the individual effects which is verified mathematically in Proposition A.2 in Appendix A.5. To efficiently generate concentration curves for multiple doses within an input window, the PK encoder uses a matrix $\textbf{w}\in\mathbb{R}^{L\times L}$ to support vectorized computation at each timestamp containing a dose, within the lag time $L$. This approach enables efficient modeling of multiple effects in $\mathcal{O}(1)$ time. Each row $t$ of w represents the elapsed time steps until the end of the window, with the time count shifting one step to the right in each subsequent row. A concentration curve is generated for each row of w using the dose at the corresponding time index. If no dose is administered at a particular time $t$, the corresponding row of w is zeroed out. The final concentration curve feature, which captures the combined impact of multiple doses, is obtained by aggregating the individual curves across the rows:
$\mathbf{w}$ $=\begin{bmatrix}0&1&2&3&\cdots&L{-}1\\ 0&0&1&2&\cdots&L{-}2\\ \vdots&\vdots&\vdots&\vdots&\ddots&\vdots\\ 0&0&0&0&\cdots&0\end{bmatrix}\in\mathbb{R}^{L\times L}$ (4)
$\widecheck{\textbf{x}}^{(i)}$ $=\sum_{t=0}^{L-1}C(\nu\textbf{w}_{t,:},\textbf{x}^{(i)}_{t},\textbf{k}^{(i)})$ (5)
Here, $\nu$ is the sampling interval of the data, $\textbf{x}^{(i)}\in\mathbb{R}^{1\times L}$ is a sparse exogenous time series treatment feature, and $\widecheck{\textbf{x}}^{(i)}\in\mathbb{R}^{1\times L}$ is the new concentration curve feature for patient $i$.
The PK encoder training algorithm 1 is described in Appendix A.6, which can be integrated into any deep learning architecture capable of handling exogenous variables in addition to modeling the target variable.
Figure 4: Our hybrid global-local architecture combines global model parameters shared across patients with patient-specific pharmacokinetic (PK) parameters. Sparse treatment time series features, such as medication doses, are input into the PK encoder to generate patient-specific treatment concentration curves $\check{\textbf{x}}$. PK encoder output $\check{\textbf{x}}$ and any other exogenous features x are used as deep learning model inputs. During each training step, the model learns relevant hidden states and output parameters $\hat{\theta}$ across patients, while $\textbf{k}^{(i)}$ values are updated for each patient $i$ using gradient descent optimization.